From mar@roe.ac.uk Fri Sep 30 16:53:38 2016 Date: Fri, 30 Sep 2016 16:53:30 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , dmr@roe.ac.uk, Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , matt.jarvis@astro.ox.ac.uk, Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 30 September 2016 WFAU VDFS Science Archives progress meeting minutes, 30 September 2016 Present: NJC, ETWS, RSC, RGM, RPB, ACD, MAR Apologies: STV, DMR, KTN, AL,MSH DoNM: 10:00 am Friday 14th October 2016 in the *** Crawford Meeting Room 1 *** ACD in the chair (room is booked). Actions discharged this time: ------------------------------------------------ ACTION: ACD to liaise with DMR and STV to establish a set of deadlines for their work. Discharged. It's been decided not to implement this. ACTION: NJC to discuss proposed VVVX meeting with RGM and to keep RSC and ACD apprised of developments pertaining to MOONS. Discharged. Actions carried forward from 9/9/2016 meeting: ---------------------------------------------- ACTION: RGM to investigate the size and content (how does it differ from DR9?) of SDSS DR10 to see if it is worth us hosting and whether it is necessary to keep all of the earlier releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. The part of this action that continues is that RGM had not yet established which early releases could safely be scrapped. RGM also noted that that DR11 and DR12 were now available and now need to be looked at. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan DR13 is installed but still need to decide which releases to keep so leaving the action for now. ACTION: MAR to look at H-ATLAS documentation. Continues. ACTION: RPB to check the log files and inform NJC of the problems with the previously listed non-survey releases. Continues ACTION: NJC to investigate and fix the problems with the previously listed non-survey releases. Continues ACTION: MSC to check with DMR about progress towards, and anticipated release date of, a usable TAP service. Continues ACTION: RPB to check with CASU whether they had any WFCAM observations made during the gap in the Edinburgh tapes. It was suggested we might like to check here first! Specific points and new actions: -------------------------------- Project management: * Nothing for the minutes WFCAM, VISTA and VST updates: * WFCAM was still off the telescope. Networking: * Nothing to report; the networks were operating nominally. WSA/VSA/OSA/GES Operations: * ETWS reported the following progress: - VPHAS reseaming is 88% done, estimated finish is 7 Oct. - WFCAM image and catalogue data is ingest up to 01/08/16 (16A complete). * RPB reported that VVV tilepawprints and ppErrBits had finished running. CU19 was being run on the current VHS release * ACD noted that the GES iDR5 linelist was available for download, ingest and release. MAR asked how the recently discovered grouting issue at CASU (data from 20150201 to 20160201) and subsequent re-processing (v1.3.1) affected the VVV release. ACTION: NJC to contact CASU, in the first instance to see what exactly is being updated, and then to discuss with Phil and Dante how this affects the current release. Hardware and Systems: * No issues with hardware to report. * RPB noted that the shuffling around of various databases to make better use of the latest servers could take place shortly, once the VHS release had finished. The recent installation of BestDR13 on ramse17 required SQL Server 2014. As all our other servers are on SQL Server 2012, for the cross mounting of databases to work, these all will need upgrading to 2014. Review of Software Release Deadlines: * Progress towards the software deadlines was reviewed briefly. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/PrepareSoft Other Software: * ETWS reported that he had: - Working on code to find re-processed v1.3.1 tiles at CASU for re-transfer. - Rewrote and streamlined the illumination correction assignment for the WSA due to the now bi-monthly creation of the tables at CASU. * NJC reported work on ingesting the Chilean deep stacks and list driven pipeline. He'd also fixed a small bug in the objId updater for combined non-surveys. Review of Survey Release Deadlines: * Progress towards forthcoming survey deadlines was reviewed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Survey Data Release: * None. The VIKING and ATLAS ESO files needed to be patched (provenance added, mag limits set correctly etc). Also the ATLAS ESO release seemed to contain more files than expected. Non-survey Data Release: * None Astrogrid deployment, VO & Data Analysis services: * Nothing to report Miscellaneous: * RGM asked about H-ATLAS. NJC replied that we had a rough idea of what we were getting at that hopefully smallish tweaks to existing curation code would see it ingested but he had not had the chance to look into it further due to other priorities. * RPB "volunteered" to book the WFAU Christmas lunch at the Queen's favourite boozer. He'd check with STV and DMR when they were likely to be in Edinburgh.